OpenStudy (vortish):

what is the x intercept and y intercept 1/5x - 1/2y =1

OpenStudy (rob1525):

when x is zero you can solve for y and when y is zero you can solve for x.

OpenStudy (vortish):

i dont understand

OpenStudy (rob1525):

so, 1/5*0-1/2y=1 then 0-1/2y=1 then solve for y by multipling by the negative reciprical so y=1*(-2/1) y=-2. do the same from x.